OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. y = 2x - 8 y = 2x - 4 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

I forgot how to set the problem up for these :/

jimthompson5910 (jim_thompson5910):

well you start with y = mx+b, your general slope intercept form m = slope b = y-intercept

jimthompson5910 (jim_thompson5910):

you are given the slope the given slope is 2, so m = 2

jimthompson5910 (jim_thompson5910):

y = mx + b y = 2x + b now you plug in the given point (3, -2)...so you replace x with 3 and y with -2 -2 = 2(3) + b -2 = 6 + b ... ... ... b = ???

OpenStudy (anonymous):

Would I subtract -2 from 6?

jimthompson5910 (jim_thompson5910):

the other way around

jimthompson5910 (jim_thompson5910):

you want to subtract 6 from both sides.

OpenStudy (anonymous):

okay, give me a moment to do that

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

http://sketchtoy.com/32604099 Like this?

jimthompson5910 (jim_thompson5910):

you got it, so m = 2 (given) b = -8 (just found) making y = mx+b turn into y = 2x - 8

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

yw